OpenStudy (anonymous):

What is the equation of a circle whose center is at the origin and whose radius is 16?

OpenStudy (anonymous):

x^2+ y^2= 256 ?

OpenStudy (campbell_st):

the general form is \[(x - h)^2 + (y - k)^2 = r^2\] centre (h, k) and radius r hope this helps you find the equation

OpenStudy (lgbasallote):

actually...you were right the first time